Ravi invests $10,000 in an investment account that pays 4% compounded semi-annually. Ravi takes each interest payment and invests it in a savings account that pays 1% compounded monthly.

a) How much money does Ravi have at the end of 10 years?

b) What is the effective annual rate he earned over 10 years?

please show full work done with all formulas
Answer #1

Semi Annual Interest Amount= $10000*4%*(6/12)=$200

Now, he will get 10*2 or 20 such payment till the end of the 10 year

Hence, semi- annual rate =(1+1%/12)^12-1=0.501%, nper(semi-annual)=10*2=20, PMT=$200, FV=?

А 1 rate 2 nper 3 pmt 4 lfv в 0.501% 20 -200 |=FV(B1,B2,B3,0,0) FV(rate, nper, pmt, [pv], [type])

ZA 1 rate 2 nper 3 pmt fv 0.501% 20 -200 $4,196.24 |

So, his total savings at the end of the 10 year=4196.24+10000=14196.24

b) His investment (PV)=$10000, FV=14196.24, nper=10, rate=?

PV FV | Η. | $(10,000.00) $14,196.24 101 =RATE(H3,0,H1,H2,0) RATE(nper, pmt, pv, [fv], [type], [guess]) nper rate

PV FV $(10,000.00) $14,196.24 10 3.57% nper rate

Hence, his effective annual rate over 10 year is 3.57%